# Date: ]June 11, 2024

### v1

@@ I Introduction
The concept of bosonic stars originated from the proposals of geons and spherical BS Wheeler (1955); Kaup (1968); RUFFINI and BONAZZOLA (1969); Bonazzola and Pacini (1966). Since then, the properties and phenomenology of these compact objects have been investigated in detail, and their formation mechanisms as well as their stability have been explored. Extensive reviews can be found, e.g., in Liebling and Palenzuela (2012); Lai (2004); Schunck and Mielke (2003). Properties of these kinds of stars strongly depend on the particular type of matter field and on a particular choice of the Lagrangian, i.e., the potential that encodes various types of self-interactions of the matter field. This allows for the modeling of various astrophysical systems, including Neutron Star (NS)-like objects, dark matter galaxy haloes, Black Hole mimickers, and intermediate-mass astrophysical objects Schunck (1998); Guzmán and Rueda-Becerril (2009); Herdeiro et al. (2021); Rosa and Rubiera-Garcia (2022); Rosa et al. (2023); Sengo et al. (2024); Adam et al. (2010). The scientific interest in the field has further increased due to the possible existence of dark-matter ultralight scalar bosons Freitas et al. (2021); Khlopov (2024) or extensions of the Standard Model, such as the axion Weinberg (1978); Wilczek (1978); Guo et al. (2023).
@@ II.2 Vectorial Boson Stars
$\textit{A}=e^{i(n\psi-wt)}\left(iVdt+\frac{H_{1}}{r}dr+H_{2}d\theta+iH_{3}\sin\theta d\psi\right),$ (12)
@@ III.2 Vectorial BS boundary conditions
$\partial_{r}F_{i}|_{r=0}=W|_{r=0}=H_{i}|_{r=0}=V|_{r=0}=0,$ (19)
@@
$\lim_{r\rightarrow\infty}F_{i}=\lim_{r\rightarrow\infty}W=\lim_{r\rightarrow\infty}H_{i}=\lim_{r\rightarrow\infty}V=0.$ (20)
@@
$\begin{split}&\partial_{\theta}F_{i}|_{\theta=0,\pi}=\partial_{\theta}W|_{\theta=0,\pi}=\partial_{\theta}H_{2,3}|_{\theta=0,\pi}=\\ &H_{1}|_{\theta=0,\pi}=V|_{\theta=0,\pi}=0.\end{split}$ (21)
@@
Figure 1: Mass frequency curves for the different models we have employed, all displayed in appendix A. Also, for different harmonic indexes. The legend shows that red dots are $n=3$ stars, greens are $n=2$, and blues are $n=1$. The different potentials for obtaining those representative curves are shown in the appendix. As can be appreciated from the plot, we have some secondary branches and also very different curves, making our set very wide-ranging.
As can be read from fig. 1, we have stars ranging from $\sim 0.2M\mu$ in the limit $w\rightarrow 1$ to $\sim 6M\mu$ for some potentials. These more massive stars were already studied in detail in Ontanon and Alcubierre (2021).
@@ IV.1 Multipole moments
For more details we refer to Adam et al. (2023); Ryan (1997c); Doneva and Pappas (2018). Once we know the functions $\nu$, $\omega$ and $B$, the coefficients are extracted straightforwardly. The combinations which provide the relevant multipoles Pappas and Sotiriou (2015) are,
@@ V Analysis
The reduced multipoles are standard definitions Yagi and Yunes (2013); Yagi et al. (2014a) and, as in Adam et al. (2022, 2023), we work with these dimensionless quantities,
@@ V.4 Comparison with Kerr multipoles and light rings
Figure 10: The upper plot illustrates the behavior of the quadrupolar moment as a function of the spin parameter across various models of PHS, clearly showing how these models converge towards the Kerr limit case of $\bar{Q}=-1$. Additionally, the plot indicates the Light Ring (LR) threshold for each model, with stars displaying LR from that threshold to the highest value in the direction of increasing $\chi$. The lower plot presents a similar analysis for the spin-octupolar moment, demonstrating the alignment between the two multipoles
@@ Appendix A Equations of interest
Table 2: Different Proca-Higgs stars models obtained varying the potential parameters.

### v2

# Date: ]August 2, 2024

@@ I Introduction
The concept of bosonic stars originated from the proposals of geons and spherical BS Wheeler (1955); Kaup (1968); RUFFINI and BONAZZOLA (1969); Bonazzola and Pacini (1966). Since then, the properties and phenomenology of these compact objects have been investigated in detail, and their formation mechanisms as well as their stability have been explored. Extensive reviews can be found, e.g., in Liebling and Palenzuela (2012); Lai (2004); Schunck and Mielke (2003). Properties of these kinds of stars strongly depend on the particular type of matter field and on a particular choice of the Lagrangian, i.e., the potential that encodes various types of self-interactions of the matter field. This allows for the modeling of various astrophysical systems, including Neutron Star (NS)-like objects, dark matter galaxy haloes, Black Hole mimickers, and intermediate-mass astrophysical objects Schunck (1998); Guzmán and Rueda-Becerril (2009); Herdeiro et al. (2021); Rosa and Rubiera-Garcia (2022); Rosa et al. (2023); Sengo et al. (2024); Adam et al. (2010). The scientific interest in the field has further increased due to the possible existence of dark-matter ultralight scalar bosons Freitas et al. (2021); Khlopov (2024) or extensions of the Standard Model, such as the axion Weinberg (1978); Wilczek (1978); Guo et al. (2023); Sakharov et al. (2021).
@@ II.2 Vectorial Boson Stars
$\textit{A}=e^{-i(n\psi+wt)}\left(iVdt+\frac{H_{1}}{r}dr+H_{2}d\theta+iH_{3}\sin\theta d\psi\right),$ (12)
@@ III.2 Vectorial BS boundary conditions
$\begin{split}&\partial_{r}\alpha|_{r=0}=\partial_{r}\beta|_{r=0}=\partial_{r}\nu|_{r=0}=\\ &W|_{r=0}=H_{i}|_{r=0}=V|_{r=0}=0,\end{split}$ (19)
@@
$\begin{split}&\lim_{r\rightarrow\infty}\alpha=\lim_{r\rightarrow\infty}\beta=\lim_{r\rightarrow\infty}\nu=\\ &\lim_{r\rightarrow\infty}W=\lim_{r\rightarrow\infty}H_{i}=\lim_{r\rightarrow\infty}V=0.\end{split}$ (20)
@@
$\begin{split}&\partial_{\theta}\alpha|_{\theta=0,\pi}=\partial_{\theta}\beta|_{\theta=0,\pi}=\partial_{\theta}\nu|_{\theta=0,\pi}=\\ &\partial_{\theta}W|_{\theta=0,\pi}=\partial_{\theta}H_{2,3}|_{\theta=0,\pi}=\\ &H_{1}|_{\theta=0,\pi}=V|_{\theta=0,\pi}=0.\end{split}$ (21)
@@
Figure 1: Mass frequency curves for the different models we use, all displayed in appendix A. The upper plot shows the solutions for harmonic index $n=1$ for the BS, PS, and PHS cases. The lower plot shows $n=2$ PS and BS and $n=3$ BS. The different potentials for obtaining those representative curves are shown in the appendix. As can be seen from the plot, we have some secondary branches and very different curves, making our set very wide-ranging.
As can be read from fig. 1, we have stars ranging from $\sim 0.2M\mu$, to $\sim 6M\mu$ for some potentials. These more massive stars were already studied in detail in Ontanon and Alcubierre (2021).
@@ IV.1 Multipole moments
For more details we refer to Adam et al. (2023); Ryan (1997c); Doneva and Pappas (2018); Pappas and Apostolatos (2012a, b). Once we know the functions $\nu$, $\omega$ and $B$, the coefficients are extracted straightforwardly. In the present study, we focus on the first few multipole moments, namely, the mass monopole $M_{0}$, spin $S_{1}$, mass quadrupole $M_{2}$, and spin octupole $S_{3}$. In terms of the metric coefficients, they are expressed as follows Pappas and Apostolatos (2012a, b)
@@
Higher order multipole moments can be evaluated using algorithms as described in Geroch (1970); Hansen (1974); Fodor et al. (1989); Backdahl and Herberthson (2005); Pappas and Apostolatos (2012a, b).
@@ V Analysis
The reduced multipoles are standard definitions commonly used throughout the literature Yagi and Yunes (2013); Yagi et al. (2014a) and, as in Adam et al. (2022, 2023), we work with these dimensionless quantities, which are completely general for any order
@@
In the above expression, $n$ may assume any integer value from 00 to $\infty$, providing a consistent formula that removes the dimensions for each multipolar order. These equations are closely connected to the multipolar expansion method, which derives the expressions found in eq. 24. While the comprehensive calculation from the metric eq. 6, through the expansion of metric functions in eq. 23, is technically elaborate, the entire procedure is detailed in Pappas and Apostolatos (2012a, b) and meticulously reviewed in section $2.2$ of Sukhov (2023).
@@ V.4 Comparison with Kerr multipoles and light rings
Figure 10: We display five PHS models, each identified by its coupling constant $\Lambda$. We present numerous solutions corresponding to the stars associated with the frequencies $w/\mu$, so the different curves are formed by the sets of distinct solutions associated with the different frequencies for each coupling constant. These models demonstrate a clear convergence to the Kerr limit case $\bar{Q}=-1$, particularly for solutions with greater compactness. Additionally, the plot indicates the Light Ring (LR) threshold for each model, with stars displaying LR from that threshold to the highest value in the direction of increasing $\chi$. The lower plot presents a similar analysis for the spin-octupolar moment, demonstrating the alignment between the two multipoles
@@ Appendix A Equations of interest
Table 2: Different Proca-Higgs stars models obtained varying the potential parameters. Recall that $v$ is related to $\alpha$ through the equation $\alpha^{2}\equiv 4\pi Gv^{2}$, as previously shown in Section III.